# MVSplat: Efficient 3D Gaussian Splatting from Sparse Multi-View Images

### v1

@@ 4.2 Main Results
Table 1: Comparison with state of the art. Running time includes both encoder and render, note that 3DGS based methods (pixelSplat and MVSplat) render dramatically faster ($\sim 500$FPS for the render). Performances are averaged over thousands of test scenes in each dataset. For each scene, the model takes two views as input and renders three novel views for evaluation. Our MVSplat performs the best in terms of all visual metrics and runs the fastest with a lightweight model size.
@@
Assessing image quality. We report results on the RealEstate10K [42] and ACID [14] benchmarks in Tab. 1. Our MVSplat surpasses all previous state-of-the-art models in terms of all metrics on visual quality, with more obvious improvements in the LPIPS metric, which is better aligned with human perception. This includes pixelNeRF [38], GPNR [26], AttnRend [6] and pixelSplat [1], with results taken directly from the pixelSplat [1] paper, and the recent state-of-the-art NeRF-based method MuRF [33], for which we re-train and evaluate its performance using the officially released code.
@@
The qualitative comparisons of the top three best models are visualized in Fig. 3. MVSplat achieves the highest quality on novel view results even under challenging conditions, such as these regions with repeated patterns (“window frames” in 1st row), or these present in only one of the input views (“stair handrail” and “lampshade” in 2nd and 3rd rows), or when depicting large-scale outdoor objects captured from distant viewpoints (“bridge” in 4th row). The baseline methods exhibit obvious artifacts for these regions, while our MVSplat shows no such artifacts due to our cost volume-based geometry representation. More evidence and detailed analysis regarding how MVSplat effectively infers the geometry structures are presented in Sec. 4.3.
Assessing model efficiency. As reported in Tab. 1, apart from attaining superior image quality, our MVSplat also shows the fastest inference time among all the compared models, accompanied by a lightweight model size, demonstrating its efficiency and practical utility. It is noteworthy that the reported time encompasses both image encoding and rendering stages. For an in-depth time comparison with pixelSplat [1], our encoder runs at 0.043s, which is more than 2×2$\times$2 × faster than pixelSplat (0.102s). Besides, pixelSplat predicts 3 Gaussians per-pixel, while our MVSplat predicts 1 single Gaussian, which also contributes to our faster rendering speed (0.0015s vs. 0.0025s) due to the threefold reduction in the number of Gaussians. More importantly, equipped with the cost volume-based encoder, our MVSplat enables fast feed-forward inference of 3D Gaussians with a much light-weight design, resulting in 10×10$\times$10 × fewer parameters and more than 2×2$\times$2 × faster speed compared to pixelSplat [1].
@@
Assessing geometry reconstruction. Our MVSplat also produces significantly higher-quality 3D Gaussian primitives compared to the latest state-of-the-art pixelSplat [1], as demonstrated in Fig. 4. pixelSplat requires an extra 50,000 steps to fine-tune the Gaussians with an additional depth regularization to achieve reasonable geometry reconstruction results. Our MVSplat instead generates high-quality geometries by training solely with photometric supervision. Fig. 4 demonstrates the feed-forward geometry reconstruction results of our MVSplat, without any extra fine-tuning. Notably, although pixelSplat showcases reasonably rendered 2D images, its underlying 3D structure contains a large amount of floating Gaussians. In contrast, our MVSplat reconstructs much higher-quality 3D Gaussians, demonstrating the effectiveness of our cost volume representation in obtaining high-quality geometry structures. To facilitate a clearer understanding of the difference between the two methods, we invite the reader to view the corresponding “.ply” files of the exported 3D Gaussians in our project page.
@@
Assessing cross-dataset generalization. MVSplat is inherently superior in generalizing to out-of-distribution novel scenes, primarily due to the fact that the cost volume captures the relative similarity between features, which remains invariant compared to the absolute scale of features. To demonstrate this advantage, we conduct two cross-dataset evaluations. Specifically, we choose models trained solely on the RealEstate10K (indoor scenes), and directly test them on ACID (outdoor scenes) and DTU (object-centric scenes). As evident in Fig. 5, MVSplat renders competitive novel views, despite scenes of the targeted datasets contain significantly different camera distributions and image appearance from those of the source dataset. In contrast, views rendered by pixelSplat degrade dramatically; the main reason is likely that pixelSplat relies on purely feature aggregations that are tied to the absolute scale of feature values, hindering its performance when it receives different image features from other datasets. Quantitative results reported in Tab. 2 further uphold this observation. Note that the MVSplat significantly outperforms pixelSplat in terms of LPIPS, and the gain is larger when the domain gap between source and target datasets becomes larger. Surprisingly, our cross-dataset generalization results on ACID even slightly surpass the pixelSplat model that is specifically trained on ACID (see Tab. 1). We attribute such results to the larger scale of the RealEstate10K training set (7$\sim$ 7$\times$ 7 × larger than ACID) and our superior generalization ability. This also suggests the potential of our method for training on more diverse and larger scale datasets in the future.
@@
Assessing more-view quality. Our MVSplat is designed to be agnostic to the number of input views, so that it can benefit from more input views if they are available in the testing phase, regardless of how many input views are used in training. We verify this by testing on DTU with 3 context views, using the model trained on the 2-view RealEstate10K dataset. Our results are PSNR: 14.30, SSIM: 0.508, LPIPS: 0.371, and pixelSplat’s are PSNR: 12.52, SSIM: 0.367, LPIPS: 0.585. Compared to the 2-view results (Tab. 2), MVSplat achieves better performance with more input views. However, pixelSplat performs slightly worse when using more views, even though we have made our best effort to extend its released 2-views-only model to support more-view testing. This suggests that the feature distribution of more views might be different from the two views one used to train pixelSplat. This discrepancy is attributed to the reliance of pixelSplat on pure feature aggregation, which lacks robustness to changes in feature distribution. This limitation is analogous to the reason why pixelSplat performs inferior in cross-dataset generalization tests discussed earlier.
@@
Table 3: Ablations on RealEstate10K. The “base + refine” is our final model, where “refine” refers to the “depth refinement” detailed in Sec. 3.1. All other abalation models are based on the “base” model by removing the mentioned module. The cost volume module plays an indispensable role in our model. Results of all models are obtained from the final converged step (300K in our experiments), except the one “w/o cross-view attention”, which suffers from over-fitting (details shown in the supplementary material Fig. A), hence we report its best performance.
Figure 6: Ablation on RealEstate10K. Each image is rendered by the corresponding ablation model indicated at its top. Colored error maps obtained by calculating the differences between the rendered images and the ground truth are attached for better comparison. All models are based on the “base” model, with the specific component being removed. “w/o cross-attn” is short for “w/o cross-view attention”. Compared to “base”, “base+refine” slightly reduces errors, “w/o cost volume” leads to the largest drop, “w/o U-Net” harms contents on the right that only exists in one input view, while “w/o cross-attn” increase the overall error intensity.
@@ 4.3 Ablations
Importance of cost volume. The cost volume serves as a cornerstone to the success of our MVSplat, which plays the most important role in our encoder to provide better geometry quality. To measure our cost volume scheme’s importance, we compare MVSplat to its variant (w/o cost volume) in Tab. 3. When removing it from the “base” model, the quantitative results drop significantly: it decreases the PSNR by more than 3dB, and increases LPIPS by 0.064 (nearly 50% relative degradation). This deterioration is more obviously evident in the rendered visual result in Fig. 6. The variant “w/o cost volume” exhibits a direct overlay of the two input views, indicating that the 3D Gaussian parameters extracted from the two input views fail to align within the same 3D space.
Ablations on cross-view attention. The cross-view matching is significant in learning multi-view geometry. The feature extraction backbone in our MVSplat is assembled with cross-view attention to enhance the feature expressiveness by fusing information between input views. We investigate it by removing the cross-view attention in our transformer. The quantitative results in Tab. 3 (“w/o cross-attn”) show a performance drop of 1dB PSNR, and the visual results in Fig. 6 (“w/o cross-attn”) showcase higher error intensity. This highlights the necessity for information flow between views. Besides, we also observe that this variant “w/o cross-attn” suffers from over-fitting (details shown in the supplementary material Fig. A), further confirming that this component is critical for model robustness. Due to the over-fitting issue, we report this variant specifically with its best performance instead of the final over-fitted one.
Ablations on the cost volume refinement U-Net. The initial cost volume might be less effective in challenging regions, thus we propose to use a U-Net for refinement. To investigate its importance, we performed a study (“w/o U-Net”) that removes the U-Net architecture. Fig. 6 reveals that, for the middle regions, the variant “w/o U-Net” render as well as the “base” mode, where content is present in both input views; but its left & right parts show obvious degraded quality in this variant compared with the “base” modes, for which content is only present in one of the inputs. This is because our cost volume cannot find any matches in these regions, leading to poorer geometry cues. In such a scenario, the U-Net refinement is important for mapping high-frequency details from input views to the Gaussian representation, resulting in an overall improvement of $\sim 0.7$ dB PSNR as reported in Tab. 3.
@@ 5 Conclusion
We present MVSplat, an efficient feed-forward 3D Gaussian Splatting model that is trained using sparse multi-view images. The key to our success is that we construct a cost volume to exploit multi-view correspondence information for better geometry learning, differs from existing approaches that resort to data-driven design. With a well-customized encoder tailored for 3D Gaussians primitives prediction, our MVSplat sets new state-of-the-art in two large-scale scene-level reconstruction benchmarks. Compared to the latest state-of-the-art method pixelSplat, our MVSplat uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization.
Limitations and Discussions. Our model might produce unreliable results for reflective surfaces like glasses and windows, which are currently open challenges for existing methods (detailed in Appendix 0.A). Besides, our model is currently trained on the RealEstate10K dataset, where its diversity is not sufficient enough to generalize robustly to in-the-wild real-world scenarios despite its large scale. It would be an interesting direction to explore our model’s scalability to larger and more diverse training datasets (e.g., by mixing several existing scene-level datasets) in the future.
@@ Appendix 0.A More Experimental Analysis
Using cost volume in pixelSplat. In the main paper, we have demonstrated the importance of our cost volume design for learning feed-forward Gaussian models. We note that such concept is general and is not specifically designed for a specific architecture. To verify this, we replace pixelSplat’s probability density-based depth prediction module with our cost volume-based approach while keeping other components intact. The results shown in Tab. A again demonstrate the importance of the cost volume by significantly outperforming the original pixelSplat, indicating the general applicability of our proposed method.
@@
Figure B: Failure cases. Our MVSplat might be less effective on the non-Lambertain and reflective surfaces.
Limitation. Our MVSplat might be less effective on non-Lambertain and reflective surfaces, as shown in Fig. B. Integrating the rendering with additional BRDF properties and training the model with more diverse datasets might be helpful for addressing this issue in the future.

### v2

@@ 4.2 Main Results
Table 1: Comparisons with the state of the art. Running time includes both encoder and render, note that 3DGS-based methods (pixelSplat and MVSplat) render dramatically faster ($\sim 500$FPS for the render). Performances are averaged over thousands of test scenes in each dataset. For each scene, the model takes two views as input and renders three novel views for evaluation. MVSplat performs the best in terms of all visual metrics and runs the fastest with a lightweight model size.
@@
Assessing image quality. We report results on the RealEstate10K [54] and ACID [21] benchmarks in Tab. 1. MVSplat surpasses all previous state-of-the-art models in terms of all metrics on visual quality, with more obvious improvements in the LPIPS metric, which is better aligned with human perception. This includes pixelNeRF [49], GPNR [35], AttnRend [10] and pixelSplat [1], with results taken directly from the pixelSplat [1] paper, and the recent state-of-the-art NeRF-based method MuRF [44], for which we re-train and evaluate its performance using the officially released code.
[1][44]
@@
The qualitative comparisons of the top three best models are visualized in Fig. 3. MVSplat achieves the highest quality on novel view results even under challenging conditions, such as these regions with repeated patterns (“window frames” in 1st row), or these present in only one of the input views (“stair handrail” and “lampshade” in 2nd and 3rd rows), or when depicting large-scale outdoor objects captured from distant viewpoints (“bridge” in 4th row). The baseline methods exhibit obvious artifacts for these regions, while MVSplat shows no such artifacts due to our cost volume-based geometry representation. More evidence and detailed analysis regarding how MVSplat effectively infers the geometry structures are presented in Sec. 4.3.
Assessing model efficiency. As reported in Tab. 1, apart from attaining superior image quality, MVSplat also shows the fastest inference time among all the compared models, accompanied by a lightweight model size, demonstrating its efficiency and practical utility. It is noteworthy that the reported time encompasses both image encoding and rendering stages. For an in-depth time comparison with pixelSplat [1], our encoder runs at 0.043s, which is more than 2×2$\times$2 × faster than pixelSplat (0.102s). Besides, pixelSplat predicts 3 Gaussians per-pixel, while our MVSplat predicts 1 single Gaussian, which also contributes to our faster rendering speed (0.0015s vs. 0.0025s) due to the threefold reduction in the number of Gaussians. More importantly, equipped with the cost volume-based encoder, our MVSplat enables fast feed-forward inference of 3D Gaussians with a much light-weight design, resulting in 10×10$\times$10 × fewer parameters and more than 2×2$\times$2 × faster speed compared to pixelSplat [1].
[1][1]
@@
Assessing geometry reconstruction. MVSplat also produces significantly higher-quality 3D Gaussian primitives compared to the latest state-of-the-art pixelSplat [1], as demonstrated in Fig. 4. pixelSplat requires an extra 50,000 steps to fine-tune the Gaussians with an additional depth regularization to achieve reasonable geometry reconstruction results. Our MVSplat instead generates high-quality geometries by training solely with photometric supervision. Fig. 4 demonstrates the feed-forward geometry reconstruction results of MVSplat, without any extra fine-tuning. Notably, although pixelSplat showcases reasonably rendered 2D images, its underlying 3D structure contains a large amount of floating Gaussians. In contrast, our MVSplat reconstructs much higher-quality 3D Gaussians, demonstrating the effectiveness of our cost volume representation in obtaining high-quality geometry structures. To facilitate a clearer understanding of the difference between the two methods, we invite the reader to view the corresponding “.ply” files of the exported 3D Gaussians on our project page.
[54][21][54][17][1][1]
@@
Assessing cross-dataset generalization. MVSplat is inherently superior in generalizing to out-of-distribution novel scenes, primarily due to the fact that the cost volume captures the relative similarity between features, which remains invariant compared to the absolute scale of features. To demonstrate this advantage, we conduct two cross-dataset evaluations. Specifically, we choose models trained solely on the RealEstate10K (indoor scenes), and directly test them on ACID (outdoor scenes) and DTU (object-centric scenes). As evident in Fig. 5, MVSplat renders competitive novel views, despite scenes of the targeted datasets containing significantly different camera distributions and image appearance from those of the source dataset. In contrast, views rendered by pixelSplat degrade dramatically; the main reason is likely that pixelSplat relies on purely feature aggregations that are tied to the absolute scale of feature values, hindering its performance when it receives different image features from other datasets. Quantitative results reported in Tab. 2 further uphold this observation. Note that the MVSplat significantly outperforms pixelSplat in terms of LPIPS, and the gain is larger when the domain gap between source and target datasets becomes larger. More surprisingly, our cross-dataset generalization results on ACID even slightly surpass the pixelSplat model that is specifically trained on ACID (see Tab. 1). We attribute such results to the larger scale of the RealEstate10K training set (7$\sim$ 7$\times$ 7 × larger than ACID) and our superior generalization ability. This also suggests the potential of our method for training on more diverse and larger-scale datasets in the future.
@@
Assessing more-view quality. MVSplat is designed to be agnostic to the number of input views, so that it can benefit from more input views if they are available in the testing phase, regardless of how many input views are used in training. We verify this by testing on DTU with 3 context views, using the model trained on the 2-view RealEstate10K dataset. Our results are PSNR: 14.30, SSIM: 0.508, LPIPS: 0.371, and pixelSplat’s are PSNR: 12.52, SSIM: 0.367, LPIPS: 0.585. Compared to the 2-view results (Tab. 2), MVSplat achieves better performance with more input views. However, pixelSplat performs slightly worse when using more views, even though we have made our best effort to extend its released 2-views-only model to support more-view testing. This suggests that the feature distribution of more views might be different from the two views used to train pixelSplat. This discrepancy is attributed to the reliance of pixelSplat on pure feature aggregation, which lacks robustness to changes in feature distribution. This limitation is analogous to the reason why pixelSplat performs inferior in cross-dataset generalization tests discussed earlier.
Figure 6: Ablations on RealEstate10K. Colored error maps obtained by calculating the differences between the rendered images and the ground truth are attached for better comparison. All models are based on the “base” model, w/o depth refinement module. “w/o cross-attn” is short for “w/o cross-view attention”. Compared to “base”, “base+refine” slightly reduces errors, “w/o cost volume” leads to the largest drop, “w/o U-Net” harms contents on the right that only exist in one input view, while “w/o cross-attn” increases the overall error intensity.
@@
Table 3: Ablations on RealEstate10K. The “base + refine” is our final model, where “refine” refers to the “depth refinement” detailed in Sec. 3.1. All other ablations are conducted on the “base” model w/o depth refinement. The cost volume module plays an indispensable role in MVSplat. Results of all models are obtained from the final converged step (300K in our experiments), except the one “w/o cross-view attention”, which suffers from over-fitting (details are shown in the supplementary material Fig. A), hence we report its best performance.
@@ 4.3 Ablations
Importance of cost volume. The cost volume serves as a cornerstone to the success of MVSplat, which plays the most important role in our encoder to provide better geometry quality. To measure our cost volume scheme’s importance, we compare MVSplat to its variant (w/o cost volume) in Tab. 3. When removing it from the “base” model, the quantitative results drop significantly: it decreases the PSNR by more than 3dB, and increases LPIPS by 0.064 (nearly 50% relative degradation). This deterioration is more obviously evident in the rendered visual result in Fig. 6. The variant “w/o cost volume” exhibits a direct overlay of the two input views, indicating that the 3D Gaussian parameters extracted from the two input views fail to align within the same 3D space.
Ablations on cross-view attention. The cross-view matching is significant in learning multi-view geometry. The feature extraction backbone in our MVSplat is assembled with cross-view attention to enhance the feature expressiveness by fusing information between input views. We investigate it by removing the cross-view attention in our transformer. The quantitative results in Tab. 3 (“w/o cross-attn”) show a performance drop of 1dB PSNR, and the visual results in Fig. 6 (“w/o cross-attn”) showcase higher error intensity. This highlights the necessity for information flow between views. Besides, we also observe that this variant “w/o cross-attn” suffers from over-fitting (details are shown in the supplementary material Fig. A), further confirming that this component is critical for model robustness. Due to the over-fitting issue, we report this variant specifically with its best performance instead of the final over-fitted one.
Ablations on the cost volume refinement U-Net. The initial cost volume might be less effective in challenging regions, thus we propose to use a U-Net for refinement. To investigate its importance, we performed a study (“w/o U-Net”) that removes the U-Net architecture. Fig. 6 reveals that, for the middle regions, the variant “w/o U-Net” render as well as the “base” model, where content is present in both input views; but its left and right parts show obvious degraded quality in this variant compared with the “base” model, for which content is only present in one of the inputs. This is because our cost volume cannot find any matches in these regions, leading to poorer geometry cues. In such a scenario, the U-Net refinement is important for mapping high-frequency details from input views to the Gaussian representation, resulting in an overall improvement of $\sim 0.7$ dB PSNR as reported in Tab. 3.
@@ 5 Conclusion
We present MVSplat, an efficient feed-forward 3D Gaussian Splatting model that is trained using sparse multi-view images. The key to our success is that we construct a cost volume to exploit multi-view correspondence information for better geometry learning, which differs from existing approaches that resort to data-driven design. With a well-customized encoder tailored for 3D Gaussians primitives prediction, our MVSplat sets new state-of-the-art in two large-scale scene-level reconstruction benchmarks. Compared to the latest state-of-the-art method pixelSplat, our MVSplat uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization.
Limitations and Discussions. Our model might produce unreliable results for reflective surfaces like glasses and windows, which are currently open challenges for existing methods (detailed in the supplementary material Appendix 0.A). Besides, our model is currently trained on the RealEstate10K dataset, where its diversity is not sufficient enough to generalize robustly to in-the-wild real-world scenarios despite its large scale. It would be an interesting direction to explore our model’s scalability to larger and more diverse training datasets (e.g., by mixing several existing scene-level datasets) in the future.
Acknowledgements This research is supported by the Monash FIT Start-up Grant. Dr. Chuanxia Zheng is supported by EPSRC SYN3D EP/Z001811/1.
@@ Appendix 0.A More Experimental Analysis
Using cost volume in pixelSplat. In the main paper, we have demonstrated the importance of our cost volume design for learning feed-forward Gaussian models. We note that such a concept is general and is not specifically designed for a specific architecture. To verify this, we replace pixelSplat’s probability density-based depth prediction module with our cost volume-based approach while keeping other components intact. The results shown in Tab. A again demonstrate the importance of the cost volume by significantly outperforming the original pixelSplat, indicating the general applicability of our proposed method.
@@
Figure B: Failure cases. Our MVSplat might be less effective on the non-Lambertian and reflective surfaces.
Limitation. Our MVSplat might be less effective on non-Lambertian and reflective surfaces, as shown in Fig. B. Integrating the rendering with additional BRDF properties and training the model with more diverse datasets might be helpful for addressing this issue in the future.